OpenStudy (anonymous):

Factor these trinomials with a POSITIVE CONSTANT ( + C) x^2+ 7x + 12 My answer: (x + 1)(x + 12 )????

OpenStudy (anonymous):

x^2+7x+12 what two numbers times to give 12 and add to give 7?

OpenStudy (anonymous):

what grade are you?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

okay so what are the factors of 12? here's one of them (1,12) what are the others?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

not quite so 1*12=12, 2*6=12, 3*4=12. So (1,12), (2,6), (3,4) are factors do any of these pairs add up to 7?

OpenStudy (anonymous):

(x + 3)(x + 4)

OpenStudy (anonymous):

perfect :D

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

x^2+ 7x + 12 = x^2 + 3x + 4x + 12 ( since 7x = 3x+4x) = x(x + 3) +4( x + 3) Take (x +3) as common term = (x +3)(x +4) Therefore, Answer = ( x+3 )( x+4 )